OpenStudy (anonymous):

A principal randomly chose 40 students from grade 5.Of these students,9 were wearing watches.Based on this information, what is the most reasonable prediction of the number of students wearing watches Ina group of 120 students from grade 5???? A.27 B. 9 C.18 D.80 Please help I'll give a medal !!!

OpenStudy (amistre64):

and your thoughts about it?

OpenStudy (amistre64):

how do we solve a proportion?

OpenStudy (amistre64):

9 out of 40 is proportionate to x out of 120 9/40 = x/120, solve for x

OpenStudy (anonymous):

..????

OpenStudy (studygurl14):

@Valerie_diaz123 You will need to cross-multiply. Do you know how to do that?

OpenStudy (anonymous):

U just do that ?? Yes I do know how to do that

OpenStudy (studygurl14):

Yes, do that. And then finish solving for x.

OpenStudy (studygurl14):

The equation that @amistre64 gave was \(\Large\frac{9}{40}=\frac{x}{120}\)

OpenStudy (anonymous):

Thanks and the answer that I got was A